# Structured Generative Modeling with The Thermodynamic Kolmogorov-Arnold Model

### v3

@@ 2.5.1 Thermodynamic Training Criterion
We inform our derivations using the Thermodynamic Integral. In place of the log-marginal likelihood outlined in Eq. 15, this exploits the following representation derived in Sec. A.4.5:
@@
The exact calculation of the integral in Eq. 28 is feasible given that the bounds are $t=0$, corresponding to the expectation being taken with respect to the prior, and $t=1$, where it is taken with respect to the posterior. However, the integral can also be expressed exactly without collapsing the temperatures, offering greater flexibility in how the integral evolves through the temperature schedule, $\{t_{k}\}_{k=0}^{N_{t}}$. In particular, the temperature schedule is discretized as:
@@ 2.5.2 Unadjusted Langevin Algorithm (ULA) and Replica Exchange
We adopt a similar strategy but follow Pang et al. (2020) in using the Unadjusted Langevin Algorithm (ULA) (Rossky et al., 1978b; Brooks et al., 2011) without MH correction. Specifically, we update samples using the forward Euler–Maruyama discretization of the Langevin SDE, as detailed by Roberts & Stramer (2002), with the transition kernel described in Sec. A.6.3:
@@
ULA is best suited for smooth, unimodal distributions, but annealing with power posteriors might improve sampling from multimodal distributions. Mixing can be further improved by periodically swapping populations between randomly selected adjacent temperatures, subject to:
@@ 3 Experiments
In Sec. 3.2, we compare the mixture EBM prior sampled using ITS against a deep EBM prior sampled using ULA. The latter shares an architectural design similar to that used by Pang et al. (2020), but with Chebyshev KAN (SS et al., 2024) energy functions. We further compare T-KAM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion. These comparisons serve to evaluate the practical strategies proposed for scaling and training T-KAM, as well as to compare ITS and our mixture EBM against ULA and deep EBMs, previously used by Pang et al. (2020).
@@
However, we recognise that quantifying the fidelity of generated data is inherently challenging and not an exact science. While the metrics used in this study have limitations, they align with those commonly found in the literature. Our conclusions are also informed by our subjective perception of the images. Furthermore, no metrics were applied in Secs. 3.1.1 &  3.1.1, as vision metrics rely on the Inception V3 model trained on ImageNet (Szegedy et al., 2015), potentially introducing bias against other datasets.
More image samples are provided in Sec. A.7.
3.1 Inductive biases enable Importance Sampling
@@ 3.1.1 MNIST & FMNIST
The MNIST (Deng, 2012) and FMNIST (Xiao et al., 2017) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq. 20. The first collection of benchmarks presented in Sec. A.2 were obtained using the hyperparameters and architecture specified for this experiment in Tab. 4. T-KAM was trained for 10 epochs with 20,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 2,000 parameter updates) Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
(a) Uniform
(b) Lognormal
(c) Gaussian
Figure 2: Exponentially-tilted priors, extracted after training on MNIST with RBFs. These plots illustrate how T-KAM can be visualized and interpreted by how its priors are adjusted relative to their reference distributions. The lognormal prior remains untrained due to posterior collapse, while the uniform prior has been skewed towards $z_{1,2}=1$. Gaussian is only slightly tilted.
(a) MNIST
(b) FMNIST
Figure 3: Generated NIST images using RBFs and different exponentially-tilted priors, initialized from $\mathcal{N}(\bm{z};\;\bm{0},\bm{1})$, $\text{Lognormal}(\bm{z};\;\bm{0},\bm{1})$, and $\mathcal{U}(\bm{z};\;\bm{0},\bm{1})$. Lognormal is inverted for clarity.
Lognormal priors resulted in posterior mode collapse. T-KAM learned to generate an average of all MNIST or FMNIST classes, and failed to learn a meaningful latent prior. In contrast, Gaussian priors produced the sharpest images for FMNIST, and uniform yielded the sharpest MNIST digits.
In Fig. 2, we show how T-KAM allows visualization of its univariate latent distributions, which is an interpretability advantage that can help us explain the observed results. Unlike Gaussian and uniform reference distributions, which provide a more symmetric and balanced initial spread over the latent space, lognormal priors introduce a strong right-skew. This asymmetric initialization makes it difficult for T-KAM to learn flexible, separable representations, ultimately impairing its ability to produce binary output features.
In contrast, the symmetric nature of Gaussian and uniform priors allows the sculpting of latent distributions and their skews to support binary generation. When cross-referencing with the priors of other latent features plotted in Sec. A.8, it becomes evident that left-skewed distributions are more successful at MNIST and FMNIST generation. This explains why the standard lognormal prior, which is right-skewed, is poorly suited to these datasets.
The experiment also demonstrates the validity of importance sampling for posterior expectations when prior-posterior alignment can be improved by choosing an appropriate reference prior, since uniform and Gaussian priors produced realistic MNIST and FMNIST samples respectively.
@@
T-KAM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=50,N_{s}=50$. The training usespanned 600 epochs with 1,000 training examples, resulting in 12,000 parameter updates. All other hyperparameters are provided in Tab. 5. Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
(a) $\mathcal{U}(\bm{z};\;\bm{0},\bm{1})$
(b) $\text{Lognormal}(\bm{z};\;\bm{0},\bm{1})$
Figure 4: Generated Darcy flow pressures with exponentially-tilted priors initialized from standard uniform and lognormal distributions. RBF bases, are contrasted against FFT bases, where RBF is colored differently for clarity.
(a) RBF prior on $z_{1,1}$
(b) FFT prior on $z_{1,1}$
Figure 5: Comparison of exponentially-tilted lognormal priors, extracted after training on Darcy between RBF bases and FFT bases. RBF basis functions led to oversmoothing and posterior collapse, while FFT bases lacked smoothness, potentially compromising adherence to the continuous assumptions imposed by KART.
Although RBFs tended to oversmooth the solution space, FFTs were more effective at capturing large-scale spatial periodicity. In cases where the prior was misspecified with a lognormal reference, RBFs failed to produce coherent samples, while FFTs still yielded coherent outputs.
This result likely stems from their suitability in representing periodic spatial relationships, a property that aligns well with the structure of the Darcy flow, as previously showcased by the success of the FNO, which also uses Fourier methods. This is somewhat reflected by Fig. 5(b), which depicts some periodicity in the learned, exponentially-tilted distribution on $z_{1,1}$. However, FFT-parameterized energy functions were also found to be non-smooth and potentially discontinuous, which is also illustrated by the distribution plotted in Fig. 5(b). This might have compromised adherence to KART and future work might explore incorporating regularization to improve their smoothness.
This shows that the choice of basis functions serves as an additional mechanism for encoding inductive bias, complementing the role of the reference prior. Once again, importance sampling proved to be viable when appropriate inductive biases were encoded.
3.2 Comparing scaling strategies against CIFAR-10 & SVHN
@@
Due to limited computational resources during this study, (detailed in Sec. A.2), training was shortened from 70,000 to 8,000 parameter updates. Given their parameter efficiency, Chebyshev bases (SS et al., 2024) were used for the prior energy functions. The deep prior architecture mirrored that of Pang et al. (2020), but it is important to note that dense KANs do not directly replicate the behavior of standard dense layers. Additionally, we used a single optimizer to train both models, rather than using separate optimizers for the prior and generator networks. We considered these discrepancies to be acceptable, given the comparative nature of this experiment.
@@
Posterior sampling was conducted with ULA, and a Gaussian reference was used, $\pi_{0}=\mathcal{N}(\bm{z};\;\bm{0},\bm{1})$. T-KAM was trained for 20 epochs with a batch size of $N_{x}=100$ on 40,000 training examples. All remaining hyperparameters are provided in Tab. 6, and architectural details for the mixture and deep prior models, as well as the CNN generators, are included in Sec. A.5.1. Annealing was conducted with $N_{t}=5$ power posteriors and the cyclic temperature scheduling outlined in Sec. A.3.1.
For brevity, only the SVHN images are visualized in the main body. Samples of CIFAR-10 are provided under Sec. A.7.5
@@
Table 1: Effectively unbiased FID and KID scores (lower is better) for SVHN.
Prior MLE SE
Mixture $\overline{\text{FID}}_{\infty}=81.02$ $\overline{\text{FID}}_{\infty}=82.13$
$\overline{\text{KID}}_{\infty}=0.0669$ $\overline{\text{KID}}_{\infty}=0.0673$
Deep $\overline{\text{FID}}_{\infty}=107.42$ $\overline{\text{FID}}_{\infty}=141.24$
$\overline{\text{KID}}_{\infty}=0.0970$ $\overline{\text{KID}}_{\infty}=0.1320$
@@
Table 2: Effectively unbiased FID and KID scores (lower is better) for CIFAR-10.
Prior MLE SE
Mixture $\overline{\text{FID}}_{\infty}=159.11$ $\overline{\text{FID}}_{\infty}=180.00$
$\overline{\text{KID}}_{\infty}=0.1600$ $\overline{\text{KID}}_{\infty}=0.1858$
Deep $\overline{\text{FID}}_{\infty}=185.10$ $\overline{\text{FID}}_{\infty}=201.75$
$\overline{\text{KID}}_{\infty}=0.1780$ $\overline{\text{KID}}_{\infty}=0.1977$
(a) SVHN with mixture prior and MLE
(b) SVHN with mixture prior and SE
Figure 6: Generated samples from T-KAM trained on SVHN using a mixture prior.
(c) SVHN with deep prior and MLE
(d) SVHN with deep prior and SE
Figure 7: Generated samples from T-KAM trained on SVHN using a deep prior.
The mixture prior with ITS outperformed the deep prior with ULA, while also being more computationally efficient and enabling faster inference, (Tab. 9). This advantage is likely due to ITS providing exact samples from the prior, whereas ULA produces only approximate, biased samples.
@@
SE did not outperform MLE and cannot be justified in this setting given its computational overhead. Alternative approaches, such as different annealing schedules or the use of autoMALA with bias correction and temperature-adaptive step size tuning, could improve performance, but they are not worth investigating until the hardware of Zetta (2024) becomes available, due to the added complexity, hyperparameter tuning, and extended training time.
The results confirm the superiority of mixture EBMs paired with ITS over deep EBMs sampled with ULA as a viable approach to scaling. However, the advantages of SE and population-based ULA remain inconclusive.
@@ 4 Conclusions
T-KAM provides a principled and scalable framework for structured and informed generative modeling. Our results demonstrate that incorporating inductive biases through both basis functions and reference priors can enable more efficient training strategies, such as importance sampling. We also show how the Kolmogorov–Arnold representation theorem can inform architectural design and introduce structural bias. Additionally, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors. While mixture priors with inverse transform sampling consistently outperformed alternative approaches across both datasets, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work. Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by Zetta (2024).
@@ 4.1.1 Hardware acceleration
Computational benchmarks are provided in Sec. A.2. The univariate and localized nature of spline interpolations complicates their parallelization on GPUs, since they involve sparse or recursive computations. Additionally, the search operations in Eqs. 1323, and 70 were parallelized on the CPU across samples, despite the remainder of T-KAM being executed on the GPU. This design stems from the fact that GPUs are generally ill-suited for parallelized search algorithms, which often rely on conditional logic. Parallelizing conditionals can lead to thread divergence, making CPUs more efficient for these operations.
While ITS enables fast inference and scales when the prior is a mixture model, T-KAM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA. Nonetheless, the unique advantages of T-KAM justify continued research and development. Notably, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control and high throughput.
The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA, (Sec. A.6.4). Even without annealing, replacing ULA with autoMALA may yield a more robust probabilistic model with stronger theoretical guarantees.
@@
4.1.2 Text-to-text, classification, and multi-modal generation
Kong et al. (2021) provided a robust framework for high-quality text-to-text generation, classification, and meta-learning, while Yuan et al. (2024) demonstrated how T-KAM might be adapted into a multi-modal generative model. This holds promise for future investigations and industry adoption, especially when the LFU becomes available. However, the symbol-vector coupling approach of Kong et al. (2021) must first be adapted to remain compatible with ITS.
4.1.3 Improved sampling for multimodal distributions
Despite its theoretical guarantees, annealing is difficult to justify given its computational cost and limited practical benefit. Future work could explore adaptations of the hierarchical EBMs proposed by Cui & Han (2024), without sacrificing the interpretability or efficiency of our model.

### v4

@@ 2.5.1 Thermodynamic Training Criterion
We inform our derivations using the Thermodynamic Integral. In place of the log-marginal likelihood outlined in Eq. 15, this uses the following representation derived in Sec. A.4.5:
@@
The exact calculation of the integral in Eq. 28 is feasible given that the bounds are $t=0$, corresponding to the expectation being taken with respect to the prior, and $t=1$, where it is taken with respect to the posterior. However, the integral can also be expressed exactly without collapsing the temperatures. This allows greater flexibility in how the integral evolves through the temperature schedule, $\{t_{k}\}_{k=0}^{N_{t}}$. In particular, the temperature schedule is discretized as:
@@ 2.5.2 Unadjusted Langevin Algorithm (ULA) and Replica Exchange
We adopt a similar strategy but follow Pang et al. (2020) in using the Unadjusted Langevin Algorithm (ULA) (Rossky et al., 1978b; Brooks et al., 2011) without MH correction. Specifically, we update samples using the forward Euler-Maruyama discretization of the Langevin SDE, as detailed by Roberts & Stramer (2002), with the transition kernel described in Sec. A.6.3:
@@
ULA is best suited for smooth, unimodal distributions, but annealing with power posteriors can improve sampling from multimodal distributions. Mixing can be further improved by periodically swapping populations between randomly selected adjacent temperatures, subject to:
@@ 3 Experiments
In Sec. 3.2, we evaluate the mixture EBM prior, sampled via component-wise ITS and paired with CNN-decoder architectures, which are smaller variants of those used by Pang et al. (2020). We additionally compare T-KAM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion. These comparisons assess the practicality of using mixture distributions for scaling T-KAM and using annealing to improve training for high-dimensional RGB images.
@@
However, we recognise that quantifying the fidelity of generated data is inherently challenging and not an exact science. While the metrics used in this study have limitations, they align with those commonly found in the literature. Our conclusions are also informed by our subjective perception of the images. Furthermore, no metrics were applied in Secs. 3.1.1 &  3.1.1, as vision metrics rely on the Inception V3 model trained on ImageNet (Szegedy et al., 2015), potentially introducing bias against other datasets, especially non-RGB images.
3.1 Importance Sampling
@@ 3.1.1 MNIST & FMNIST
The MNIST (Deng, 2012) and FMNIST (Xiao et al., 2017) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq. 20. The first collection of benchmarks presented in Sec. A.2 were obtained using the hyperparameters and architecture specified for this experiment in Tab. 4. T-KAM was trained for 10 epochs with 50,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 5,000 parameter updates). Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
(a) Real samples
(b) No ref. prior, FFT basis
(c) Gaussian ref. prior, RBF basis
Figure 2: Generated MNIST examples using different bases and exponentially-tilted priors
(a) Real samples
(b) No ref. prior, RBF basis
(c) Gaussian ref. prior, RBF basis
Figure 3: Generated FMNIST examples using different bases and exponentially-tilted priors
T-KAM generated a diverse set of samples when a non-restrictive reference prior was used, (Gaussian or no reference), thus validating importance sampling and independent energy-based priors for simple datasets. Subjectively, the best MNIST samples arose with no reference prior and RBF bases, whereas the best FMNIST samples came from a Gaussian reference prior with RBF bases, potentially because FMNIST has smoother features than MNIST.
The additional samples plotted in Sec. A.7.1 illustrate that the choice and combination of bases and reference priors are important. Poorly chosen combinations led to posterior collapse and low-quality samples, thereby showcasing the effect of inductive bias in guiding sample quality.
In Sec. A.8, examples of trained univariate latent distributions have been plotted to demonstrate T-KAM’s interpretability. Learning can be understood by how T-KAM reshaped its latent representation relative to the initial reference prior. For instance, lognormal priors showed little change, indicating posterior collapse, and FFTs caused much sharper relationships than the smoother RBFs.
@@
T-KAM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=100,N_{s}=100$, and the first likelihood model in Eq. 20. Training spanned 1,000 epochs with 1,000 training examples, resulting in 10,000 parameter updates. All other hyperparameters are provided in Tab. 5. Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
(a) Real samples
(b) Gaussian ref. prior, FFT basis
(c) Uniform ref. prior, RBF basis
Figure 4: Generated Darcy flow examples using different bases and exponentially-tilted priors
RBFs tended to oversmooth the solution space while FFTs were more effective at capturing large-scale spatial periodicity. This likely stems from their suitability in representing periodic spatial relationships, a property that aligns well with the structure of the Darcy flow, as previously showcased by the success of the FNO, which also uses Fourier methods. This is somewhat reflected by Fig. 32, which depicts some periodicity in the learned, exponentially-tilted distributions on $z_{q,p}$.
However, FFT-parameterized energy functions were also found to be non-smooth and potentially discontinuous, which is also illustrated by the distributions plotted in Fig. 32. This might have compromised adherence to KART and future work might explore incorporating regularization to improve their smoothness. Overall, the results show that the choice of basis functions serves as an additional mechanism for encoding inductive bias, complementing the role of the reference prior.
Once again, importance sampling only proved to be viable when appropriate inductive biases were encoded, and some combinations of bases and reference priors failed to learn meaningful features, (demonstrated by the extra samples of Sec. A.7.3).
3.2 Langevin Monte Carlo and Steppingstone estimator
@@
We evaluated T-KAM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015). The first likelihood model in Eq. 20 was adopted for both datasets. Due to limited computational resources during this study, (detailed in Sec. A.2), training was shortened to 8,000 parameter updates, (70,000 were used in the study by Pang et al. (2020)), and smaller versions of the CNN decoders used by Pang et al. (2020) were adopted as the generator architectures.
Additionally, a single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator networks. Chebyshev KAN bases (SS et al., 2024) were used for the prior energy functions, and full architectural details are provided under Sec. A.5.1.
@@
Posterior sampling was conducted with ULA, and the standard normal reference prior was used, $\pi_{0}=\mathcal{N}\left(\bm{z};\;\bm{0},\bm{1}\right)$. T-KAM was trained for 40 epochs with a batch size of $N_{x}=100$ on 20,000 training examples. All remaining hyperparameters are provided in Tab. 6. Annealing was conducted with $N_{t}=10$ power posteriors and the cyclic temperature scheduling outlined in Sec. A.3.1.
@@
Table 1: Effectively unbiased FID and KID scores (lower is better) for SVHN ($32\times 32$).
MLE SE
$\overline{\text{FID}}_{\infty}=97.90$ $\overline{\text{FID}}_{\infty}=145.64$
$\overline{\text{KID}}_{\infty}=0.0827$ $\overline{\text{KID}}_{\infty}=0.1407$
@@
Table 2: Effectively unbiased FID and KID scores (lower is better) for CelebA ($64\times 64$).
MLE SE
$\overline{\text{FID}}_{\infty}=226.88$ $\overline{\text{FID}}_{\infty}=209.30$
$\overline{\text{KID}}_{\infty}=0.2590$ $\overline{\text{KID}}_{\infty}=0.2138$
(a) Real samples
(b) Maximum Likelihood
(c) Steppingstone Estimator
Figure 5: Generated SVHN ($32\times 32$) examples using mixture prior architecture with inverse transform sampling from the prior and Langevin sampling from the posterior
(a) Real samples
(b) Maximum Likelihood
(c) Steppingstone Estimator
Figure 6: Generated CelebA ($64\times 64$) examples using mixture prior architecture with inverse transform sampling from the prior and Langevin sampling from the posterior
Despite being trained with reduced complexity for a small iteration count, our mixture prior model was able to generate low quality samples of SVHN and CelebA. This demonstrates the potential of component-wise ITS for exact and highly efficient sampling in replacement to the approach of Pang et al. (2020), which relied on ULA to sample from the prior, thus introducing bias and inefficiency.
@@
SE only resulted in better image quality for CelebA and cannot be justified in this setting given its computational overhead. Alternative approaches, such as different annealing schedules or the use of autoMALA with bias correction and temperature-adaptive step size tuning, could improve performance, but they are not worth investigating until the hardware of Zetta (2024) becomes available, due to the added complexity, hyperparameter tuning, and extended training time.
@@ 4 Conclusions
T-KAM provides a principled and scalable framework for structured and informed generative modeling. Our results demonstrate that correctly chosen basis functions and reference priors can enable more efficient training on simple or well-understood datasets through importance sampling. We also show how the Kolmogorov-Arnold representation theorem can inform architectural design and introduce structural bias, allowing T-KAM’s architecture to be guided entirely by the size of its latent space. Moreover, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors in terms of speed and exact sampling. However, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work. Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by Zetta (2024).
@@ 4.1.1 Hardware acceleration
Benchmarks are provided in Sec. A.2. The univariate, localized nature of spline interpolations complicates their parallelization on GPUs, since they involve sparse and recursive computations. The search operations in Eqs. 1323, and 69 were also difficult to parallelize on GPUs, which are generally ill-suited for search algorithms, as they often rely on conditional logic. Parallelizing conditionals can lead to thread divergence, making GPUs inefficient for these operations.
Additionally, while ITS enables fast inference and scales when the prior is a mixture model, T-KAM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA. GPUs are not well suited to Langevin Monte Carlo methods due to the irregular memory access patterns and bandwidth pressure caused by differentiating with respect to inputs, which requires storing the outputs of each layer within the generator.
Nonetheless, the unique advantages of T-KAM justify continued research and development. Moreover, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control that can mitigate the challenges of sparsity, recursion, and irregular memory access patterns. The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA, (Sec. A.6.4). Even without annealing, replacing ULA with autoMALA may yield a more robust probabilistic model with stronger theoretical guarantees.
@@
4.1.2 More opportunities for importance sampling
For SVHN and CelebA, IS proved infeasible as a posterior sampler, so ULA was used for Sec. 3.2. This may be attributed to prior–posterior mismatch, which rendered the IS estimator ineffective. Moreover, relying on IS restricts the prior to its initial domain, since proposals are generated from the prior using ITS and finite-domain KAN energy functions. In contrast, ULA can explore values of $\bm{z}$ beyond this domain, enabling adaptation of the prior domain within the grid-updating scheme for KANs proposed by Liu et al. (2024).
A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe & Cadima, 2016), similar to the approach of Kingma et al. (2014), who also applied PCA to train VAEs. To address prior–posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins. Additionally, efforts should be directed toward selecting an appropriate initial domain or adapting the domain dynamically during training with IS.
4.1.3 Expressivity
The requirements for ITS outlined in Sec. 2.1 are quite restrictive, which may limit expressivity at scale. Additionally, it is often desirable to enrich the latent space with more expressive distributions. For example, Normalizing Flows (Papamakarios et al., 2021) achieve this by applying a sequence of reversible transformations that progressively map a simple base distribution into a complex one.
This might be achievable in T-KAM by introducing more latent distributions, or revising Eq. 12 into a Mixture of Experts framework (Cai et al., 2024) where $\alpha_{q,p}$ could be realized as a more flexible kernel to enable different inductive biases. However, this might be expensive for the model by Pang et al. (2020) since its use of Langevin sampling is computationally intense. For example, the previous work of Cui & Han (2024), which divides their latent EBM space using the diffusion strategy by Ho et al. (2020), states that sampling is expensive due to their use of MCMC methods.
T-KAM could potentially circumvent the additional cost by using IS as a fast sampler, though this would require addressing its current limitations.